OpenStudy (anonymous):

12(2x - 8) = 4(5 - 2x) + 3(7 - 3x)

OpenStudy (anonymous):

its a fraction

OpenStudy (anonymous):

12(2x-8)=24x-96 4(5-2x)+3(7-3x)=20-8x+21-9x=41-17x 24x-96=41-17x Add 96 to 41 and you get 24x=137-17x Add 17x to 24x and you get 41x=137 Then divide 41/137 and you get x=0.29927